OpenStudy (anonymous):

State the parent function of the equation y = 1+3square root of x+2 and the transformation that were applied

OpenStudy (anonymous):

\[y = 1+3\sqrt{x+2}\]

OpenStudy (anonymous):

@cwrw238 Think you could help me out? :)

OpenStudy (anonymous):

@jim_thompson5910 Can you help?

jimthompson5910 (jim_thompson5910):

Start with y = sqrt(x) Then shift 2 units to the left to get y = sqrt(x+2) Then stretch it vertically by a factor of 3 ----> y = 3*sqrt(x+2) Then shift 1 unit up ----> y = 1 + 3*sqrt(x+2)

OpenStudy (anonymous):

I'm lost at the first part. :/ What's sqrt?

jimthompson5910 (jim_thompson5910):

sqrt(x) means \(\Large \sqrt{x}\) "square root of x"

jimthompson5910 (jim_thompson5910):

it's short hand for writing "square root of " everytime

OpenStudy (anonymous):

Haha, ok.